# WINTER on S250206dm: A near-infrared search for an electromagnetic counterpart

### v1

@@ Abstract
We present near-infrared follow-up observations of the International Gravitational Wave Network (IGWN) event S250206dm with the Wide-Field Infrared Transient Explorer (WINTER). WINTER is a near-infrared time-domain survey designed for electromagnetic follow-up of gravitational-wave sources localized to 300 deg2. The instrument’s wide field of view (1.2 deg2), dedicated 1-m robotic telescope, and near-infrared coverage (0.9-1.7 microns) are optimized for searching for kilonovae, which are expected to exhibit a relatively long-lived near-infrared component. S250206dm is the only neutron star merger in the fourth observing run (to date) localized to 300,deg2 with a False Alarm Rate below one per year. It has a probability of being a neutron star–black hole (NSBH) merger and a probability of being a binary neutron star (BNS) merger, with a credible region spanning 38 deg2, an estimated distance of 373 Mpc, and an overall false alarm rate of approximately one in 25 years. WINTER covered of the probability area at least once and at least three times. Through automated and human candidate vetting, all transient candidates found in WINTER coverage were rejected as kilonova candidates. Unsurprisingly, given the large estimated distance of 373 Mpc, the WINTER upper limits do not constrain kilonova models. This study highlights the promise of systematic infrared searches and the need for future wider and deeper infrared surveys.
1 Introduction
@@
The discovery of GW170817 during the second Advanced LIGO-Virgo observing run (O2) marked a breakthrough in multi-messenger astronomy, providing the first confirmed detection of a kilonova powered by r-process nucleosynthesis in the neutron-rich ejecta of a binary neutron star (BNS) merger (Abbott et al., 2017a, c, b). This event established the importance of joint observations of gravitational wave (GW) alerts and identifying electromagnetic (EM) counterparts and investigating the physics of compact object mergers, the neutron star equation of state, and heavy-element production (e.g., Abbott et al. 2017b; Barnes et al. 2016; Coulter et al. 2017; Barnes & Kasen 2013; Evans et al. 2017; Goldstein et al. 2017; Grossman et al. 2014; Haggard et al. 2017; Hallinan et al. 2017; Kasen et al. 2017, 2013; Kasliwal et al. 2019; Li & Paczynski 1998; Margutti et al. 2017; Metzger et al. 2010; Metzger 2020; Roberts et al. 2011; Rosswog 2005; Tanaka & Hotokezaka 2013; Tanvir et al. 2017; Troja et al. 2017).
@@
Despite extensive optical and infrared follow-up campaigns, subsequent LIGO-Virgo observing runs (e.g., the third observing run, O3) yielded no confirmed EM counterparts to BNS or neutron star-black hole (NSBH) mergers (e.g., Coughlin et al., 2020; Kasliwal et al., 2020; Abbott et al., 2021a, b; Collaboration et al., 2021; Zhu et al., 2021a). Many of these events were predicted to be too faint for detection by optical telescopes, as kilonovae exhibit viewing-angle-dependent blue emission that fades rapidly (1 week), while a longer-lived near-infrared (NIR) signal is predicted to be more isotropic (Kasen et al., 2013, 2017; Metzger, 2020). Indeed, GW170817 had significantly longer-lived infrared emission compared to optical emission (Tanvir et al., 2017). Zhu et al. (2021b) predicts kilonovae may be up to 8–10 times more detectable in the near-infrared compared to optical wavelengths. However, systematic sensitive wide-field infrared searches have been limited by high detector costs and high sky backgrounds. Some shallow searches were performed by the Palomar Gattini-IR survey (Moore et al., 2016; De et al., 2020) during O3 (Coughlin et al., 2019). Two new wide-field IR surveyors have come online in the fourth observing run, including the Wide-field Infrared Transient Explorer (WINTER), discussed in this study, and PRIME a new NIR survey on a 1.8-meter telescope with four large-format HgCdTe sensors covering a 1.56 deg2 field of view (Durbak et al., 2024; Kutyrev et al., 2023; Yama et al., 2023).
@@
In this paper, we present results from WINTER follow-up of the LIGO–Virgo–KAGRA (LVK; Aasi et al. 2015; Acernese et al. 2014; Somiya 2012; Aso et al. 2013; Akutsu et al. 2020) gravitational wave trigger S250206dm (LIGO–Virgo–KAGRA Collaboration, 2025a). S250206dm was first detected by the LVK at 2025-02-06T21:25:30.439 UT, as part of the ongoing O4 run. The event was recovered by the GW search pipelines GstLAL (Messick et al., 2017; Sachdev et al., 2019; Hanna et al., 2020; Cannon et al., 2021; Ewing et al., 2024; Tsukada et al., 2023), MBTA (Adams et al., 2016; Aubin et al., 2021) and PyCBC Live (Allen et al., 2012; Allen, 2005; Dal Canton et al., 2020; Usman et al., 2016; Nitz et al., 2017; Davies et al., 2020), with an estimated distance of 373 Mpc and a false alarm rate of about 1 in 25 years. The event, as reported by the LVK, has a 55% probability of being a NSBH merger, with a 37% probability of being a BNS merger (LIGO–Virgo–KAGRA Collaboration, 2025a); in addition, the probability that the lighter compact object is a neutron star is (LIGO–Virgo–KAGRA Collaboration, 2025b). S250206dm’s localization underwent several revisions, with the latest (released 36 hours after merger) having a 50% localization area of 38 deg2, and a 90% area of 547 deg2. With its relatively precise localization and confident detection, S250206dm is one of the most promising merger candidates in O4 with a high probability of containing a neutron star.
In addition to the gravitational wave properties of S250206dm which suggested a possible electromagnetic counterpart, the IceCube Neutrino Observatory (Aartsen et al., 2017) promptly reported the detection of two neutrinos in spatial and temporal coincidence with the event (IceCube Collaboration, 2025), as part of its realtime search for neutrino-gravitational wave coincident detections (see e.g Abbasi et al., 2023). Neutrino 1 was detected 109s before merger, while neutrino B was detected 289s after merger.111https://roc.icecube.wisc.edu/public/LvkNuTrackSearch/output/2025_02_06_S250206dm-7-Update.html We performed additional ToO observations of neutrino fields, similar to other teams (Becerra et al., 2025; Masi, 2025).
In this paper, we report on the nondetection of an EM counterpart to S250206dm, observed with WINTER during O4. We describe our observational strategy, data processing, and implications for future NIR follow-up of GW events.
2 WINTER follow-up campaign
2.1 WINTER observatory
@@
The WINTER telescope is operated by the WINTER Supervisory Program (WSP), a multithreaded python-based control system that manages scheduling and observations using PyQt5 for interfacing with the Qt framework. Daemon processes on the observatory’s control PC initiate WSP, which automates key observatory functions, including nightly scheduling, target-of-opportunity (ToO) handling, and priority-based command execution, while allowing manual overrides when necessary. Drawing on methodologies from MINERVA (Swift et al., 2015), Robo-AO (Riddle et al., 2012), and GOTO (Dyer et al., 2018), WSP facilitates automated decision-making for telescope operations, including dome control and observation execution. Target of Opportunity (ToO) observations are submitted automatically through a dedicated API system, which validates scheduling requests and sends them to the scheduler.
2.2 S250206dm observations
@@
WINTER observations of the GW localization region commenced on UT 2025-02-08T04:05:52, 30.5 hours after the merger. No observations were possible on the first night after the merger due to poor weather at Palomar Observatory, which continued to limit observations throughout the follow-up campaign. WINTER observations are conducted on a predefined field grid. An optimal schedule to tile the GW skymap using WINTER fields was computed using gwemopt (Coughlin et al., 2018, 2019; Dietrich et al., 2020). Each WINTER field was observed in the J-band with a set of eight two-minute long dithered exposures, corresponding to a combined integration time of sixteen minutes per field. These fields were subsequently reobserved on the nights of UT 2025 Feb 9, 10, 11, 16, and 17. All WINTER pointings overlaid on the GW-skymap are shown in Figure 1. The full details of the pointings are available on TreasureMap (Wyatt et al., 2020). Within ten days of merger, a total of 43% of the probability enclosed within the latest available GW skymap at the time of writing222https://gracedb.ligo.org/api/superevents/S250206dm/files/Bilby.offline1.multiorder.fits was observed with WINTER at least once, 40% of the total probability was observed at least twice, and 35% was observed at least three times.
2.3 Data processing and transient detection
@@
The WINTER images were processed by the WINTER data reduction pipeline implemented in mirar333https://github.com/winter-telescope/mirar (Karambelkar & Stein et al., in prep) – a modular, python-based framework designed for real-time, end-to-end processing of astronomical images to search for transients. A detailed description of the processing steps will be provided in a forthcoming publication. Briefly, the pipeline begins by performing dark subtraction, flat calibration and sky-subtraction on the raw WINTER images. Astrometric solutions are then computed using the packages Astrometry.net (Lang et al., 2010) and Scamp (Bertin, 2006). Eight astrometry-calibrated dithers are then stacked using the software Swarp (Bertin et al., 2002), and the stacked images are photometrically calibrated relative to the Two Micron All-Sky Survey (2MASS; Skrutskie et al., 2006) point source catalog (Cutri et al., 2003). Images from each of WINTER’s six detectors are reduced separately, so each sixteen-minute observation of a field produces six stacked images each with dimensions 0.6 deg x 0.3 deg. The median depths achieved during our observations of S250602dm for the six detectors were 17.6, 17.6, 17.4, 18.0, 17.5, and 17.4 mag (AB) respectively. Of the total GW skymap probability, 41% was observed to a depth of 17 mag or deeper, 38% was observed to a depth of 17.5 mag or deeper, 31% was observed to a depth of 18 mag or deeper, and 11% was observed to a depth of 18.5 mag or deeper.
@@ 2.4 Candidate Searches
2.4.1 Candidates With Multiple WINTER detections
@@
2.4.2 Candidates with a ZTF crossmatch
@@
2.4.3 Candidates with a TNS cross-match
We select all candidates included on the TNS event page for S250206dm, which was 116 at the time of writing.444https://www.wis-tns.org/system/files/ligo/o4/S250206dm_20250206_212530/S250206dm_20250206_212530_AFTER.json. Of these, 95 lie within the 95% contour, though just four TNS transients fell within fields observed by WINTER. We crossmatch all WINTER candidates to the TNS transients with a radius of 3", but find no coincident detections.
@@ 2.4.4 Candidate Vetting
We crossmatch all sources to PS1 to a star/galaxy morphology classifier trained using PS1 data (Tachibana & Miller, 2018). We reject WINTER sources within 7" of bright (m) stars, and those within 3" of fainter stars.
@@
Bogus sources with poor subtraction residuals, especially those detections close to poorly-subtracted reference image sources. We require candidates to have at least two detections which pass visual inspection and appear to be plausibly real.
@@
3 Comparison with kilonova models
We generate model light curves for kilonovae using NSBH and BNS models presented in Ahumada et al., in prep. These models are simulated with the latest version (Bulla, 2023) of the 3D Monte Carlo radiative transfer code possis (Bulla, 2019) assuming a two-component (dynamical+wind) ejecta configuration and using state-of-the-art heating rates (Rosswog & Korobkin, 2024) and opacities (Tanaka et al., 2020). The kilonova grids span the expected parameter space, with the NSBH grid varying the average velocity of the dynamical ejecta and the masses of the dynamical and wind ejecta (for a total of 108 models) and the BNS grid the mass, average velocity and average electron fraction for both components (for a total of 3072 models). To account for distance uncertainties, we sample 100 distances from the probability distribution function for S250206dm, as provided by the multi-order Bilby skymap (LIGO–Virgo–KAGRA Collaboration, 2025b). For each sampled distance, we compute a full model grid of calculated ejecta properties to explore the range of possible emission scenarios.
Figure 2 presents the resulting marginalized model prediction, obtained by interpolating all light curves onto a common time grid and computing summary statistics at each epoch. We present the -band median light curve and confidence intervals and compare to the -band median light curve to demonstrate the relatively bright and persistent NIR plateau. Additionally, we overlay the limiting magnitudes of all WINTER follow-up observations (as described in Section 2) to assess whether the models are observationally constrained.

### v2

@@ Abstract
We present near-infrared follow-up observations of the International Gravitational Wave Network (IGWN) event S250206dm with the Wide-Field Infrared Transient Explorer (WINTER). Near-infrared observations are a critical component of electromagnetic follow-up to gravitational-wave events, as kilonovae are expected to exhibit long-lived emission at these wavelengths, especially from lanthanide-rich ejecta. WINTER is a near-infrared time-domain survey facility designed for EM follow-up of gravitational-wave sources, featuring a wide field of view (1.2 deg2), a dedicated 1-m robotic telescope, and coverage spanning 0.9–1.7 microns. S250206dm is the only neutron star merger in the fourth observing run, to date, localized to with a False Alarm Rate below one per year, making it a particularly valuable target for follow-up. It has a 55% probability of being a neutron star–black hole (NSBH) merger and a 37% probability of being a binary neutron star (BNS) merger. The event’s estimated distance is 373 Mpc, with a 50% credible region spanning 38 deg2. WINTER covered 43% of the probability area at least once and 35% at least three times. Through automated and human candidate vetting, all transients were rejected as kilonova candidates. Given the large distance of the event, the WINTER upper limits do not place meaningful constraints on kilonova models. However, similar observations of future events—or in combination with optical surveys—can begin to exclude portions of the kilonova model space. This study highlights the promise of systematic infrared searches and the need for future wider and deeper infrared surveys.
I Introduction
@@
The discovery of GW170817 during the second Advanced LIGO–Virgo observing run (O2) marked a breakthrough in multi-messenger astronomy, providing the first confirmed detection of a kilonova powered by r-process nucleosynthesis in the neutron-rich ejecta of a binary neutron star (BNS) merger (Abbott et al., 2017b, c, a). This event demonstrated the power of coordinated gravitational wave (GW) and electromagnetic (EM) observations for understanding the astrophysics of compact object mergers.
Subsequent analyses of GW170817 placed constraints on the neutron star equation of state (e.g., Margalit and Metzger, 2017; Coughlin et al., 2018; Breschi et al., 2021), confirmed the origin of short gamma-ray bursts (e.g., Goldstein et al., 2017; Troja et al., 2017), and provided compelling evidence that neutron star mergers contribute significantly to heavy-element production (e.g., Li and Paczynski, 1998; Roberts et al., 2011; Rosswog, 2005; Metzger et al., 2010; Barnes and Kasen, 2013; Kasen et al., 2013; Grossman et al., 2014; Tanaka and Hotokezaka, 2013; Barnes et al., 2016; Kasen et al., 2017; Metzger, 2020). Observational efforts during and after the event provided unprecedented EM coverage across the spectrum (e.g., Coulter et al., 2017; Evans et al., 2017; Haggard et al., 2017; Hallinan et al., 2017; Margutti et al., 2017; Kasliwal et al., 2019; Tanvir et al., 2017).
While dozens of GW candidate events with a neutron star have been detected since GW170817, few have been both promising in their classification and well localized enough to enable meaningful EM follow-up. S250206dm, the focus of this study, is notable as the only neutron star merger candidate in the fourth observing run (O4) to date with a 90 localization area smaller than 600 deg2 and a false alarm rate below one per year (LIGO–Virgo–KAGRA Collaboration, 2025a; Pillas et al., 2025). With a probability that one component is a neutron star and a relatively compact skymap released within 36 hours, S250206dm presented one of the most compelling targets for infrared follow-up during O4.
@@
Despite extensive optical and infrared follow-up campaigns, LIGO-Virgo observing runs throughout O3 and O4a yielded no confirmed EM counterparts to BNS or neutron star-black hole (NSBH) mergers (e.g., Coughlin et al., 2020; Kasliwal et al., 2020; Abbott et al., 2021a, b; Collaboration et al., 2021; Zhu et al., 2021b; Pillas et al., 2025). Many of these events were predicted to be too faint for detection by optical telescopes, as kilonovae exhibit viewing-angle-dependent blue emission that fades rapidly (1 week) (Kasen et al., 2013, 2017; Metzger, 2020). In contrast, the near-infrared (NIR) emission can persist for over a week, with a slower decline timescale and reduced angular dependence due to lanthanide-rich ejecta. This longer duration makes the NIR signal more accessible for follow-up, particularly for events with delayed or large localization. GW170817, for example, had significantly longer-lived infrared emission compared to optical emission (Tanvir et al., 2017). Models such as Zhu et al. (2021a) predicts kilonovae may be up to 8–10 times more detectable in the near-infrared compared to optical wavelengths.
However, systematic sensitive wide-field infrared searches have been limited by high detector costs and high sky backgrounds. Some shallow searches were performed by the Palomar Gattini-IR survey (Moore et al., 2016; De et al., 2020) during O3 (Coughlin et al., 2019a). Two new wide-field IR surveyors have come online in the fourth observing run, including the Wide-field Infrared Transient Explorer (WINTER), discussed in this study, and PRIME a new NIR survey on a 1.8-meter telescope with four large-format HgCdTe sensors covering a 1.56 deg2 field of view (Durbak et al., 2024; Kutyrev et al., 2023; Yama et al., 2023). WINTER and PRIME, located in the northern and southern hemispheres respectively, offer complementary sky coverage and observing strategies. For example, both followed up the O4a NSBH candidate S240422ed, each covering a substantial portion of the localization region with only partial overlap near its center111https://treasuremap.space/alerts?graceids=S240422ed.
@@
This study presents results from WINTER’s follow-up of the gravitational-wave event S250206dm, detected by the LIGO–Virgo–KAGRA (LVK) collaboration (Aasi et al., 2015; Acernese et al., 2014; Somiya, 2012; Aso et al., 2013; Akutsu et al., 2020). The trigger occurred on 2025 February 6 at 21:25:30.439 UT and was publicly reported via the Gamma-ray Coordinates Network (GCN) (LIGO–Virgo–KAGRA Collaboration, 2025a).
The event was identified independently by three low-latency GW search pipelines: GstLAL (Messick and others, 2017; Sachdev and others, 2019; Hanna and others, 2020; Cannon et al., 2021; Ewing et al., 2024; Tsukada and others, 2023), MBTA (Adams et al., 2016; Aubin and others, 2021), and PyCBC Live (Allen, 2005; Allen et al., 2012; Usman et al., 2016; Nitz et al., 2017; Dal Canton et al., 2020; Davies et al., 2020). Based on these analyses, S250206dm had an estimated luminosity distance of 373 Mpc and a false alarm rate of roughly 1 in 25 years. According to the LVK’s initial classification, the event has a 55% probability of being a NSBH merger and a 37% probability of being a BNS merger (LIGO–Virgo–KAGRA Collaboration, 2025a). Notably, the posterior probability that the lighter compact object is a neutron star exceeds 99% (LIGO–Virgo–KAGRA Collaboration, 2025b). S250206dm’s sky localization was refined over time, with the final skymap—released 36 hours after the merger—yielding a 50% credible region of 38 deg2 and a 90% region of 547 deg2. This combination of confident astrophysical classification and relatively compact localization makes S250206dm one of the most promising neutron star merger candidates of the O4 observing run.
In addition to the gravitational wave properties of S250206dm which suggested a possible electromagnetic counterpart, the IceCube Neutrino Observatory (Aartsen et al., 2017) promptly reported the detection of two neutrinos in spatial and temporal coincidence with the event (IceCube Collaboration, 2025), as part of its realtime search for neutrino-gravitational wave coincident detections (see e.g Abbasi et al., 2023). Neutrino 1 was detected 109s before merger, while neutrino B was detected 289s after merger.222https://roc.icecube.wisc.edu/public/LvkNuTrackSearch/output/2025_02_06_S250206dm-7-Update.html We performed additional ToO observations of neutrino fields, similar to other teams (Becerra et al., 2025; Masi, 2025).
In this paper, we report on the non-detection of an EM counterpart to S250206dm, observed with WINTER during O4. We describe our observational strategy, data processing, and implications for future NIR follow-up of GW events.
II WINTER follow-up campaign
II.1 WINTER observatory
@@
WINTER is designed for autonomous, low-latency response to transient alerts, with the ability to begin observations within 10 minutes of trigger receipt. The telescope is operated by the WINTER Supervisory Program (WSP), a multithreaded python-based control system that handles nightly scheduling, dome operations, and real-time prioritization of Target of Opportunity (ToO) events. ToO requests are submitted through a dedicated API and processed automatically without the need for human intervention.
II.2 S250206dm observations
@@
WINTER observations of the GW localization region commenced on UT 2025-02-08T04:05:52, 30.5 hours after the merger. No observations were possible on the first night after the merger due to poor weather at Palomar Observatory, which continued to limit observations throughout the follow-up campaign. WINTER observations are conducted on a predefined field grid. An optimal schedule to tile the GW skymap using WINTER fields was computed using gwemopt (Coughlin et al., 2018, 2019b; Dietrich et al., 2020). Each WINTER field was observed in the J-band with a set of eight two-minute long dithered exposures, corresponding to a combined integration time of sixteen minutes per field. These fields were subsequently reobserved on the nights of UT 2025 Feb 9, 10, 11, 16, and 17. All WINTER pointings overlaid on the GW-skymap are shown in Figure 1. The full details of the pointings are available on TreasureMap (Wyatt et al., 2020). Within ten days of merger, a total of 43% of the probability enclosed within the latest available GW skymap at the time of writing333https://gracedb.ligo.org/api/superevents/S250206dm/files/Bilby.offline1.multiorder.fits was observed with WINTER at least once, 40% of the total probability was observed at least twice, and 35% was observed at least three times.
II.3 Data processing and transient detection
@@
The WINTER images were processed by the WINTER data reduction pipeline implemented in mirar444https://github.com/winter-telescope/mirar (Karambelkar & Stein et al., in prep) – a modular, python-based framework designed for real-time, end-to-end processing of astronomical images to search for transients. A detailed description of the processing steps will be provided in a forthcoming publication555See https://mirar.readthedocs.io/en/latest/autogen/winter.default.html for a flowchart of all steps involved.. Briefly, the pipeline begins by performing dark subtraction, flat calibration and sky-subtraction on the raw WINTER images. Astrometric solutions are then computed using the packages Astrometry.net (Lang et al., 2010) and Scamp (Bertin, 2006). Eight astrometry-calibrated dithers are then stacked using the software Swarp (Bertin et al., 2002), and the stacked images are photometrically calibrated relative to the Two Micron All-Sky Survey (2MASS; Skrutskie et al., 2006) point source catalog (Cutri et al., 2003). Images from each of WINTER’s six detectors are reduced separately, so each sixteen-minute observation of a field produces six stacked images each with dimensions 0.6 deg x 0.3 deg. The median depths achieved during our observations of S250602dm for the six detectors were 17.6, 17.6, 17.4, 18.0, 17.5, and 17.4 mag (AB) respectively. Of the total GW skymap probability, 41% was observed to a depth of 17 mag or deeper, 38% was observed to a depth of 17.5 mag or deeper, 31% was observed to a depth of 18 mag or deeper, and 11% was observed to a depth of 18.5 mag or deeper.
@@ II.4 Candidate searches
II.4.1 Candidates with multiple WINTER detections
@@
II.4.2 Candidates with a ZTF crossmatch
@@
II.4.3 Candidates with a TNS cross-match
We also check against the Transient Name Server, which serves as the official community repository of all astrophysical transients. We select all candidates reported to TNS which lie within the S250206dm skymap (accessed via the dedicated TNS event page for S250206dm). There were 116 at the time of writing.666https://www.wis-tns.org/system/files/ligo/o4/S250206dm_20250206_212530/S250206dm_20250206_212530_AFTER.json. Of these, 95 lie within the 95% contour, though just four TNS transients fell within fields observed by WINTER. We crossmatch all WINTER candidates to the TNS transients with a radius of 3", but find no coincident detections.
@@ II.4.4 Candidate Vetting
We crossmatch all sources to PS1 to a star/galaxy morphology classifier trained using PS1 data (Tachibana and Miller, 2018). We reject WINTER sources within 7" of bright (m) stars, and those within 3" of fainter stars.
@@
Bogus sources with poor subtraction residuals, especially those detections close to poorly-subtracted reference image sources. We require candidates to have at least two detections which pass visual inspection and appear to be plausibly real. A handful of candidates which nominally had two detections were rejected at this stage after careful visual comparison to UKIRT reference images and other sources within the field, on the grounds that at least one of the apparent detections was spurious.
@@
III Comparison with kilonova models
We generate model light curves for kilonovae using NSBH and BNS models presented in Ahumada et al., in prep. These models are simulated with the latest version (Bulla, 2023) of the 3D Monte Carlo radiative transfer code possis (Bulla, 2019) assuming a two-component (dynamical+wind) ejecta configuration and using state-of-the-art heating rates (Rosswog and Korobkin, 2024) and opacities (Tanaka et al., 2020). The kilonova grids span the expected parameter space, with the NSBH grid varying the average velocity of the dynamical ejecta and the masses of the dynamical and wind ejecta (for a total of 108 models) and the BNS grid the mass, average velocity and average electron fraction for both components (for a total of 3072 models). To account for distance uncertainties, we sample 100 distances from the probability distribution function for S250206dm, as provided by the multi-order Bilby skymap (LIGO–Virgo–KAGRA Collaboration, 2025b), resulting in a distribution with a mean of 375 Mpc and a 1 (68) width of 107 Mpc. For each sampled distance, we compute a full model grid of calculated ejecta properties to explore the range of possible emission scenarios.
Figure 2 presents the combined model prediction, obtained by interpolating all light curves onto a common time grid and computing summary statistics at each epoch. We present the -band median light curve and confidence intervals and compare to the -band median light curve to demonstrate the relatively bright and persistent NIR plateau. Additionally, we overlay the limiting magnitudes of all WINTER follow-up observations (as described in Section II) to assess whether the models are observationally constrained.